A thermal neutron with a speed v corresponding to the average thermal energy at
temperature T = 300 K is incident on a crystal. Will a diffraction pattern be
observed? Explain.
1
Expert's answer
2019-03-07T11:29:30-0500
"\dfrac{1}{2}mv^2=E=k_B T \Rightarrow v=\sqrt{\dfrac{2k_B T}{m}}"

"\lambda=\dfrac{h}{mv}=\dfrac{h}{\sqrt{2k_B Tm}}"

"\lambda=\dfrac{6.626 \cdot 10^{-34}}{\sqrt{2\cdot 300 \cdot 1.38 \cdot 10^{-23} \cdot 1.674 \cdot 10^{-27}}} \approx 1.78 \cdot 10^{-10}\, m"

If the wavelength "\lambda" is on the same order as the atomic separations, then neutrons will diffract when passing through the crystal.